OpenStudy (anonymous):

if f(x) = 3x -7, then f^-1 (x)=

OpenStudy (anonymous):

um what now?

OpenStudy (anonymous):

any help?

OpenStudy (anonymous):

i dont understand what you said

OpenStudy (anonymous):

no... yeah i know wow must be stupid

OpenStudy (anonymous):

lazy no i want to understand it i have a big test coming up am try to understand it not get the answer

OpenStudy (dmndlife24):

Still need help

OpenStudy (dmndlife24):

@wesley2514 to find the inverse of a function, you have to first change the "f(x)" to a "y"...like this y = 3x - 7 The second step is to interchange the x and y variables... x = 3y - 7 The third step is to just solve for y... x + 7 = 3y x + 7/3 = y Your last step is to plug the "f^-1(x)" back in for y, like this... f^-1 (x) = \[\frac{ x+7 }{ 3 }\]